### v1

# LLM-TabFlow: Synthetic Tabular Data Generation with Inter-column Logical Relationship Preservation

Abstract
Synthetic tabular data have widespread applications in industrial domains such as healthcare, finance, and supply chains, owing to their potential to protect privacy and mitigate data scarcity. However, generating realistic synthetic tabular data while preserving inter-column logical relationships remains a significant challenge for the existing generative models. To address these challenges, we propose LLM-TabFlow, a novel approach that leverages Large Language Model (LLM) reasoning to capture complex inter-column relationships and compress tabular data, while using Score-based Diffusion to model the distribution of the compressed data in latent space. Additionally, we introduce an evaluation framework, which is absent in literature, to fairly assess the performance of synthetic tabular data generation methods in real-world contexts. Using this framework, we conduct extensive experiments on two real-world industrial datasets, evaluating LLM-TabFlow against other five baseline methods, including SMOTE (an interpolation-based approach) and other state-of-the-art generative models. Our results show that LLM-TabFlow outperforms all baselines, fully preserving inter-column relationships while achieving the best balance between data fidelity, utility, and privacy. This study is the first to explicitly address inter-column relationship preservation in synthetic tabular data generation, offering new insights for developing more realistic and reliable tabular data generation methods.
keywords:
Tabular Data Generation, Inter-Column Relationship, Large Language Models, Diffusion Models, Data Reasoning, Generative AI, Synthetic Data
1 Introduction
Synthetic data is artificially generated using original data and a model trained to replicate the characteristics and structure of the original dataset (Jordon et al., 2022). It is widely recognized as a solution for enabling secure and privacy-preserved data sharing while also addressing challenges associated with limited access to real data (Long et al., 2025a). Given the critical need for privacy-preserved data exchange and the scarcity of real data, synthetic data, particularly synthetic tabular data, has the potential for widespread application in industrial scenarios, such as supply chain analytics, operations optimization, and digital twin simulation.
Tabular data, often structured in a table with rows and columns and stored in relational databases (Vieira et al., 2020), is ubiquitous (van Breugel and van der Schaar, 2024) across various industries—ranging from medical records in healthcare and transaction data in finance to sales data in retail to inventory data in warehousing. Such data is often carefully designed and maintained by domain experts, with data schemes or structures that are domain-specific and where columns are interdependent. Compared to unstructured data modalities such as image, video, and text, tabular data is structurally simpler yet semantically more complex, featuring well-defined columns, rich inter-column interactions, high dimensionality and heterogeneity, as well as domain-specific context and complexity. Thus, synthetic tabular data generation is challenging and remains underexplored, with its development significantly lags behind these unstructured data modalities (van Breugel and van der Schaar, 2024).
Nevertheless, many methods (Chawla et al., 2002; Xu et al., 2019; Kotelnikov et al., 2023; Borisov et al., 2022)—mainly adapted from models designed for generating other data modalities—have been proposed in literature to generate tabular data. One of earliest such approaches is the Synthetic Minority Over-sampling TEchnique (SMOTE), originally proposed by Chawla et al. (2002) to address dataset imbalance problem. As a data augmentation method, this interpolation-based method has recently been found effective in generating tabular data (Wang et al., 2024a). However, despite its promising performance on several benchmark datasets, as reported in Long et al. (2025b), SMOTE inherently struggles to preserve data privacy effectively and capture complex column dependencies. To overcome these limitations, deep generative models—leveraging their success in image, text, and audio synthesis—have been increasingly explored for tabular data generation. For instance, Xu et al. (2019) proposed CTGAN, a generative adversarial network (GAN)-based method specifically designed for tabular data. CTGAN learns the distribution of tabular data and generates synthetic rows from the learned representation. Although CTGAN improves privacy preservation (Zhao et al., 2024), it struggles to capture the interdependencies between categorical and continuous columns (Zhang et al., 2023). More recently, diffusion model-based methods such as TabDDPM (Kotelnikov et al., 2023) and TabSyn (Zhang et al., 2023) have been introduced to synthesize tabular data by learning the denoising process in latent space. These approaches demonstrate superior performance compared to CTGAN, particularly in preserving intricate tabular structures. Furthermore, Borisov et al. (2022) proposed GReaT, a large language model (LLM)-based approach for tabular data generation. GReaT transforms tabular data into textual representations and fine-tunes generative models to produce synthetic data, enabling the learning of underlying distributions without requiring complex encoding and decoding processes.
While these generative methods demonstrate effectiveness in generating tabular data for machine learning tasks—particularly for classification and prediction—they often overlook or fail to explicitly preserve inter-column relationship in the generated data, a property crucial for real-world applications. Inter-column relationships are often thus governed by domain-specific constraints (Long et al., 2025b), such as the temporal requirement that delivery times must occur after ordering times in logistics or the mathematical constraint that the product of unit price and quantity must equal the total amount in procurement. Due to their failure to preserve inter-column relationships, existing generative methods—primarily optimized for facilitating machine learning tasks—may produce unrealistic tabular data that lack real-world utility, limiting their applicability in simulation and informed decision-making. Moreover, despite the critical importance of inter-column relationships in industrial tabular data generation, few prior studies have explicitly focused on this aspect.
In this paper, we thus aim to bridge this gap. First, we propose an evaluation framework to assess the effectiveness of synthetic tabular data in real-world scenarios. Next, we introduce LLM-TabFlow, a LLM-based synthetic tabular data generation method designed to preserve both privacy and inter-column relationships. This approach leverages LLMs to compress tabular data by inferring relationships and employs diffusion models to train and sample synthetic data in the latent space. Finally, we conduct experiments to comprehensively evaluate the proposed method and compare it with both traditional and state-of-the-art approaches. Experimental results on two real-world industrial datasets demonstrate that LLM-TabFlow outperforms existing methods under the proposed evaluation framework across three key dimensions: data fidelity, data utility, and data privacy, achieving a better trade-off between data fidelity and privacy preservation. By effectively preserving the complex inter-column dependencies, LLM-TabFlow is particularly well-suited for real-world applications beyond traditional machine learning tasks. It is the first method, to the best of our knowledge, to explicitly focus on capturing complex inter-column relationships in tabular data generation. By prioritizing relational consistency within tables, LLM-TabFlow sets a new benchmark for creating realistic and reliable synthetic datasets.
The rest of this paper is structured as follows. Section 2 reviews related work. Section 3 details the proposed approach, LLM-TabFlow. Section 4 presents experimental settings, including datasets, baselines, along with the evaluation framework. Section 5 presents experimental results. Section 6 discusses the limitation and implications of the proposed approach. Finally, Section 7 concludes the paper and summarizes future work.
2 Related Work
This section reviews related work, including the preliminaries, the complexities of tabular data, and relevant methods for generating synthetic tabular data.
2.1 Tabular Data and Its Inherent Complexity
Tabular data, typically structured in a table with rows and columns and stored in relational databases, is the dominant modality (Vieira et al., 2020; van Breugel and van der Schaar, 2024) in many fields, such as healthcare (Nik et al., 2023; Hernandez et al., 2022), finance (Sattarov et al., 2023), and supply chains (Long et al., 2025a). While modern generative models have achieved human-like performance in synthesizing other data modalities, such as text, images, and audio, as demonstrated by ChatGPT (Wu et al., 2023), DALL-E 2 (Lin, 2023), and Suno (Yu et al., 2024), the development of synthetic tabular data remains in its early stage and significantly lags behind these advancements.
This disparity is largely due to the complexity inherent in tabular data, which differs fundamentally from other data modalities. Unlike unstructured data modalities—such as text and images—that are typically one- or two-dimensional and follow well-defined spatial and temporal distributions, tabular data, though structured, is more complex and often lacks clear structural patterns, featuring both categorical and continuous columns with rich inter-column dependencies. Moreover, the lack of a unified format across features, each governed by its own distinct distribution, further complicates the generation of high-quality synthetic tabular data (Shan et al., 2019; Wang et al., 2024b). This heterogeneity makes it challenging to capture and preserve the underlying inter-column relationships columns while ensuring data fidelity and utility. These relationships (Long et al., 2025b), whether mathematical, temporally sequential, or geographically hierarchical, are often defined based on domain-specific knowledge and differ significantly from one domain to another. As a result, although effective for other data modalities, current state-of-the-art methods aimed at enhancing machine learning tasks struggle to preserve inter-column relationships—a critical aspect for ensuring the utility of synthetic tabular data in real-world applications. This highlightthe need for more focused research in this aspect.
2.2 Evaluation Framework
Currently, most research on tabular data generation methods is validated using low-dimensional datasets, such as Adult Income (Kohavi et al., 1996) and California Housing (Pace and Barry, 1997). In contrast, fewer studies explore large-scale datasets like HIGGS (Baldi et al., 2014) and Covertype (Asuncion et al., 2007). However, even these large-scale datasets predominantly consist of numerical attributes and lack complex inter-column relationships, particularly between categorical and numerical data. This limits their ability to represent real-world scenarios where logical consistency and dependencies are critical. And recent advancements in synthetic tabular data generation have been evaluated on real-world industrial datasets, particularly in finance and healthcare. For example, studies have used datasets such as Payments and Credit (Sattarov et al., 2023) and MIMIC-III (Johnson et al., 2016). These datasets inherently involve complex inter-column relationships, such as transaction amounts linked to account balances in financial data or patient diagnoses influencing treatment plans in healthcare. Despite the presence of these relationships, evaluations in these studies have largely overlooked their preservation. Instead, the focus has been on traditional metrics like machine learning utility and statistical alignment (Papadaki et al., 2024; van Breugel et al., 2024). Specifically, the current evaluation of data fidelity primarily focuses on capturing pairwise column correlations and multivariate dependencies (Margeloiu et al., 2024). While these metrics provide insights into statistical relationships, they fail to guarantee the consistency or preservation of logical relationships within multivariate distributions (Margeloiu et al., 2024). For example, in industrial tabular data from supply chains, complex inter-column relationships are often present, such as order dates being logically linked to delivery dates or total sales prices calculated as a function of price, quantity, and discounts (Suh et al., 2024; Borisov et al., 2022). Moreover, some logical relationships such as hierarchical consistency (e.g., city and country) are not unified. And conditional dependencies between those features may change depending on the specific combination of values in each row. This variability adds another layer of complexity to accurately modeling and evaluating these relationships. Recognizing this challenge, (Long et al., 2025b) proposed the first evaluation framework specifically designed for inter-column relationship preservation. Their work highlights the need for advanced generative models capable of accurately modeling these intricate, row-specific dependencies.
2.3 Tabular Data Generation Approaches
One of the earliest effective methods for generating tabular data is the Synthetic Minority Over-sampling TEchnique (SMOTE), originally proposed by Chawla et al. (2002) to address dataset imbalance. As a data augmentation method, this interpolation-based method has recently been found effective in generating tabular data (Wang et al., 2024a). Despite its simplicity, SMOTE achieves performance comparable to many advanced generative models on several benchmark datasets, as reported by Long et al. (2025b). However, since SMOTE generates new data points by linearly interpolating among neighboring data points without modeling the underlying data distribution, it inherently struggles to preserve data privacy effectively and is less capable of capturing complex relationships and dependencies between variables.
Therefore, deep generative models have been introduced for tabular data generation. Xu et al. (2019) proposed TVAE, a Variational Autoencoder-based method, for generating tabular data. They further presented Conditional Generative Adversarial Networks (CTGAN), which achieved better performance than TVAE in terms of data accuracy. CTGAN transforms input categorical data into numerical formats using techniques such as one-hot encoding (Kim et al., 2022) and analog bits encoding (Zheng and Charoenphakdee, 2022). However, these encoding techniques often struggle to capture the rich dependencies and interactions in the input (Lee et al., 2023). This limitation restricts CTGAN from effectively recovering the complex logical relationships in real-world tabular data.
Inspired by the remarkable success of diffusion models in in image (Ho et al., 2020) and audio generation (Huang et al., 2023), diffusion model-based methods have been proposed for tabular data generation, such as TabDDPM (Kotelnikov et al., 2023) and TabSyn (Zhang et al., 2023). Unlike TVAE and CTGAN, which rely on adversarial training and may struggle with issues like mode collapse or imbalanced data, diffusion model-based methods use a gradual denoising process, ensuring smoother and more stable learning of the tabular data distribution. These methods have shown to outperform TVAE and CTGAN in terms of better preserving inter-column relationships. However, the denoising process, an essential component of diffusion-model-based methods, is less effective for tabular data due to its heterogeneity, unlike image data, which is a more uniform in nature (Xu et al., 2019). Thus, diffusion models, with their isotropic noise structure, also struggles to effectively capture the logical relationships among columns (Long et al., 2025b).
More recently, the effectiveness of pretrained language models (PLMs) in natural language processing has led to the emergence of a new class of tabular data generation methods—LLM-based tabular data generation methods, such as GReaT (Borisov et al., 2022), REaLTabFormer (Solatorio and Dupriez, 2023), and Tabula (Zhao et al., 2023). These methods convert tabular data into tokenized formats before fine-tuning PLMs, enabling them to generate synthetic data that preserve the structure and semantics of the original dataset. However, compared to GAN-based and diffusion model-based methods, they require significantly longer fine-tuning and data generation times. Moreover, because PLMs rely heavily on sequential modeling, they are restricted their ability to model complex multivariate relationships when dependencies are not sequential. Therefore, the class of methods, GReaT, often fail to capture critical inter-column dependencies, reducing the fidelity of synthetic data. methods like GReaT often fail to capture critical inter-column dependencies, reducing the fidelity of synthetic data. While graph-based approaches, such as PAFT (Xu et al., 2024d), improve dependency preservation, they depend heavily on prior knowledge, which may not always be available or accurate.
While existing methods demonstrate effectiveness in privacy preservation and in maintaining sequential inter-column relationships to some extent, they fall short in accurately preserving more complex inter-column dependencies in the generated synthetic data. We thus address this gap in the paper, proposing a LLM-based tabular data generation approach—LLM-TabFlow—which will be detailed in the following section.
3 The LLM-TabFlow Approach
Recently, LLMs such as ChatGPT (Brown et al., 2020), LLaMa (Touvron et al., 2023), and Deepseek (Guo et al., 2025) have been demonstrated remarkable success in natural language processing and beyond. These models enable the automation of complex data processing tasks and significantly enhance the semantic understanding of text, including tabular data. For instance, Li et al. (2024) demonstrated their ability to summarize tabular data from reports, while Liu et al. (2023) showcased their potential for textual and symbolic reasoning on tables, facilitating deeper analysis and complex query resolution. These advancements highlight LLMs’ ability to bridge natural language understanding and structured data analysis. However, pretraining LLMs on large volumes of tabular text data (Hegselmann et al., 2023) demands significant computational resources and time. Additionally, tabular data lacks the structured context LLMs rely on, making them less effective for generation tasks compared to their strengths in text understanding. In contrast, deep generative models, such as diffusion models, excel at capturing complex patterns in tabular data by learning latent representations. These models efficiently model the joint distribution of features, enabling high-quality synthetic data generation. While deep generative models are adept at replicating statistical correlations, LLMs offer unique advantages in understanding semantic relationships and preserving inter-column dependencies.
Therefore, we propose LLM-TabFlow, which integrate LLMs reasoning into the deep generative models for tabular data generation. This method leverages deep generative models for latent space exploration while utilizing LLMs for semantic reasoning, ensuring both statistical accuray and domain-consistent relationships in the generated data. The workflow of LLM-TabFlow is illustrated in Figure 1, with each step detailed in the following sections.
Figure 1: Overview of the workflow of the LLM-TabFlow approach.
3.1 Serialization
The first step of LLM-TabFlow is serialization, as shown in Figure 1. Let $X=[x_{1},x_{2},\dots,x_{n}]$ represent the tabular dataset, where $x_{i}$ corresponds to the values of the $i$-th column, and $C=[c_{1},c_{2},\dots,c_{n}]$ denote the corresponding column names. Additionally, let $D=[d_{1},d_{2},\dots,d_{n}]$ denotes the column descriptions, where $d_{i}$ describes the semantics or metadata of the column $c_{i}$; $Y$ represents the target data (dependent variable). In this step, the column names and descriptions are serialized into natural language using the following mapping function:
$F:(c_{j},d_{j})\mapsto c_{j}\text{ : }d_{j},j=1,2,\dots,n$ (1)
where $c_{j}$ represents a column name and $d_{j}$ its corresponding description. Once the dat is serialized, the next task is to use the serialized data, $F(c_{j},d_{j})$ as input to a LLM to infer logical groupings among columns. The relationships between the columns are determined based on these serialized descriptions. The inferred relationships, combined with expert-designed domain knowledge, is used to create a general prompt $p$ and task-specific instructions $i$, forming the input for the LLM. This input is represented as $(F(c_{j},d_{j}),p,i)$, where the LLM utilizes this information to identify meaningful relationships among multiple columns. Finally, the LLM generates lists of column names that are logically connected based on the serialized data and instructions provided.
3.2 LLM Reasoning
The second step is LLM reasoning. We illustrate examples of prompts and instruction inputs for key logical relationships in tabular data in Figure 1, including hierarchical consistency and logical dependencies, such as mathematical and temporal relationships. The output of LLM reasoning is structured as dictionaries and lists to represent inter-column relationships. Specifically, hierarchical consistency is represented as a set of mappings:
=$\mathcal{R}^{\mathcal{H}}=\Bigl{\{}c^{h}_{1,j}:\{c^{h}_{2,j},c^{h}_{3,j},\dots,c^{h}_{n,j}\}\Bigr{\}},$ (2)
where $c^{h}_{1,j}$ is the most granular column in the semantic hierarchy of the $j$ th row, and $\{c^{h}_{2,j},c^{h}_{3,j},\dots,c^{h}_{n,j}\}$ are other columns within the same hierarchical group $h$. Here, $h$ can also take on values such as $1,2,3,\ldots,k$, representing different hierarchical groups. Besides, $n$ represents the number of additional columns that belong to the same hierarchical group. Mathematical dependencies are defined as a mapping between independent and derived columns:
=$\mathcal{R}^{\mathcal{M}}=\Bigl{\{}\{f^{m}_{i_{1}},f^{m}_{i_{2}},\dots,f^{m}_{i_{n}}\}:\{f^{m}_{d_{1}},f^{m}_{d_{2}},\dots,f^{m}_{d_{q}}\}\Bigr{\}},$ (3)
where $\{f^{m}_{i_{1}},f^{m}_{i_{2}},\dots,f^{m}_{i_{n}}\}$ are independent values of n columns, and $\{f^{m}_{d_{1}},f^{m}_{d_{2}},\dots,f^{m}_{d_{q}}\}$ are derived columns that depend on them. The values of $q$ and $n$ indicate the number of columns in the respective sets. Additionally, $t\in\{1,2,3,\ldots,k\}$, where $m$ indicates the mathematical dependency group. Temporal dependencies are represented as an ordered sequence of columns:
$\mathcal{R}^{\mathcal{T}}=\Bigl{[}f^{t}_{1},f^{t}_{2},\dots,f^{2}_{n}\Bigr{]}$ (4)
where $f^{t}_{1},f^{t}_{2},\dots,f^{2}_{n}$ are columns arranged in chronological order, ensuring that temporal relationships between features are maintained. Additionally, $t\in\{1,2,3,\ldots,k\}$, where $t$ indicates the dependency group. The parameter $n$ denotes the total number of columns in the ordered sequence.
3.3 Tabular Data Compression
The relationships $\mathcal{R}^{\mathcal{H}}$, $\mathcal{R}^{\mathcal{M}}$, and $\mathcal{R}^{\mathcal{T}}$ are preserved through the LLMs reasoning and integral to compressing complex tabular data, simplifying its structure and accelerating the subsequent data generation process. Specifically, $\mathcal{R}^{\mathcal{H}}$ is applied to identify and preserve columns at the most granular hierarchical level (e.g., “City”), while columns at higher levels in the hierarchy (e.g., “Region” and “Country”) are removed. Similarly, $\mathcal{R}^{\mathcal{M}}$ identifies independent columns within mathematical relationships and retains only those, discarding derived columns to reduce the complexity of the data structure. For temporal relationships, $\mathcal{R}^{\mathcal{T}}$ is used to retain the column with the earliest time point while transforming other columns within the temporal group $\{f^{t}_{1},f^{t}_{2},\dots,f^{2}_{n}\}$. The transformation involves computing the differences between consecutive time points, denoted as $dt_{k}=f^{t}_{k}-f^{t}_{k-1},\quad\text{for }k=2,3,\dots,p$, where $d^{t}_{k}$ represents the difference between the $k$-th time column and the $(k-1)$-th time column. This transformation simplifies the tabular data representation to be $\mathcal{X}_{\text{c}}$.
3.4 Compressed Data Generation
Figure 2: Illustration of compressed data generation via score-based diffusion in the latent space.
The compressed data $\mathcal{X}_{\text{c}}$ is used to train a score-based diffusion model for generating synthetic tabular data $\mathcal{X}_{\text{c}}^{{}^{\prime}}$. Following the preprocessing methods in Zhang et al. (2023), we integrate a tokenizer and a Transformer-based Variational Autoencoder (VAE) into a unified encoder. Specifically, we first tokenize $\mathcal{X}_{\text{c}}$ to represent mixed data types (numerical and categorical) within a unified dense vector space. The tokenized data is then mapped into a latent space via the Transformer VAE, where the encoder outputs the mean and log variance of the latent distribution. By reparameterizing the latent distribution, we derive the latent embedding, denoted as $\mathbf{H}_{0}$, which serves as the initial representation of the data in the latent space. The diffusion model is then employed to learn the underlying distribution of $\mathbf{H}_{0}$ by progressively transforming it through a forward process and subsequently recovering it via a reverse process. To model the latent distribution $p(\mathbf{H}_{0})$, we employ a score-based diffusion framework (Karras et al., 2022), which consists of a forward stochastic perturbation process and a reverse generative process. The forward process incrementally introduces Gaussian noise to the initial latent embedding $\mathbf{H}_{0}$, transforming it into a noisy latent state $\mathbf{H}_{T}$ over time. This evolution is described by the following equation:
$\mathbf{H}_{t}=\mathbf{H}_{0}+\alpha(t)\boldsymbol{\eta},\quad\boldsymbol{\eta}\sim\mathcal{N}(0,\mathbf{I})$ (5)
where $\alpha(t)$ represents the noise scaling function at time $t$, determining the intensity of the perturbation applied to the latent representation. The reverse process seeks to reconstruct the original latent state by iteratively denoising $\mathbf{H}_{t}$. This is governed by the following stochastic differential equation:
$d\mathbf{H}_{t}=-2\alpha^{\prime}(t)\alpha(t)\nabla_{\mathbf{H}_{t}}\log p(\mathbf{H}_{t})\,dt+\sqrt{2\alpha^{\prime}(t)\alpha(t)}\,d\mathbf{W}_{t}$ (6)
where $\nabla_{\mathbf{H}_{t}}\log p(\mathbf{H}_{t})$ is the learned score function, which estimates the gradient of the log-density and guides the denoising process. And $\alpha^{\prime}(t)$ represents the time derivative of the noise scaling function $\alpha(t)$, meaning it quantifies how the noise level changes over time. The term $\mathbf{W}_{t}$ denotes the standard Wiener process, introducing controlled stochasticity into the reverse dynamics. To achieve accurate reconstruction, the model is trained using denoising score matching. Specifically, a neural network $\boldsymbol{\eta}_{\phi}(\mathbf{H}_{t},t)$ is optimized to approximate the noise added during the forward process. The training objective is to minimize the following loss function (Karras et al., 2022):
$\mathcal{L}=\mathbb{E}_{\mathbf{H}_{0}\sim p(\mathbf{H}_{0})}\mathbb{E}_{t\sim p(t)}\mathbb{E}_{\boldsymbol{\eta}\sim\mathcal{N}(0,\mathbf{I})}\|\boldsymbol{\eta}_{\phi}(\mathbf{H}_{t},t)-\boldsymbol{\eta}\|^{2}_{2}.$ (7)
Once the diffusion model is trained, synthetic latent embeddings $\mathbf{H}^{\prime}$ can be sampled by applying the learned reverse process. These embeddings are subsequently decoded to reconstruct synthetic data representations that preserve the statistical and structural properties of the original dataset. This generative mechanism ensures that the synthetic data $\mathcal{X}_{\text{c}}^{\prime}$ remains faithful to the distribution of the original compressed data $\mathcal{X}_{\text{c}}$, facilitating high-quality synthetic data generation.
3.5 Synthetic Data Decompression
The compressed synthetic data $\mathcal{X}_{\text{c}}^{{}^{\prime}}$ is decompressed into a reconstructed dataset $\mathcal{X}^{{}^{\prime}}$ using relationship-guided mappings $\mathcal{R}_{h}$, $\mathcal{R}_{m}$, and $\mathcal{R}_{t}$, which encode hierarchical, mathematical, and temporal relationships, respectively. The reconstructed dataset $\mathcal{X}^{{}^{\prime}}$ is obtained as:
$\mathcal{X}^{{}^{\prime}}=f_{\text{decom}}\bigl{(}\mathcal{X}_{\text{c}}^{{}^{\prime}},\mathcal{R}_{h},\mathcal{R}_{m},\mathcal{R}_{t}\bigr{)},$ (8)
where $f_{\text{decom}}:\mathcal{X}_{\text{c}}^{{}^{\prime}}\times\mathcal{R}_{h}\times\mathcal{R}_{m}\times\mathcal{R}_{t}\rightarrow\mathcal{X}^{{}^{\prime}}$ is the decompression function. Here, $\mathcal{X}^{{}^{\prime}}$ represents the dataset that faithfully reconstructs the original data structure, while preserving the hierarchical, mathematical, and temporal relationships encoded by $\mathcal{R}_{h}$, $\mathcal{R}_{m}$, and $\mathcal{R}_{t}$.
4 Experimental Settings
This section presents the experimental setup used to assess the performance of the proposed approach, as well as the state-of-the-art tabular data generation methods for comparison. It includes details on the evaluation datasets, the framework used for evaluation, and the hyperparameter configurations.
4.1 Datasets
Table 1: Statistics of the two evaluation datasets.
Dataset
# of
Rows
# of
Num. Feat.
# of
Cat. Feat.
Size of
Training Set
Size of
Test Set
Retail 172,765 26 15 155,488 17,277
Purchasing 29,590 14 7 26,631 2,959
Two real-world industrial datasets were adopted for experimental evaluation. The first dataset is an open-source dataset [^1] https://data.mendeley.com/datasets/8gx2fvg2k6/3 from the retail sector, denoted as Retail in this paper. The second dataset, denoted as Purchasing, is a proprietary dataset on procurement that comes from a manufacturing company. The statistics of these two datasets are presented in Table 1.
As shown in Table 1, the Retail dataset, with over 150,000 samples, captures the complex interactions between suppliers, customers, and logistics, integrating geographical, temporal, financial, and product-related information. It includes 26 numerical and 15 categorical columns, modeling transactional activities and the relationships between variables such as payment type, profit per order, and sales per customer. The dataset’s complexity arises from the interplay between elements like product categories, customer segments, and order details interact across various levels, influencing sales, pricing, and profitability. Additionally, geographic data, such as customer and order locations, links regional sales patterns to product performance. Similarly, the Purchasing dataset reflects the complex dynamics of procurement, with temporal, categorical, and numerical features. It models order cycles, lead times, and stakeholder interactions, with financial attributes like prices, currencies, and discounts adding additional layer of complexity by capturing transactional values and tax implications. This combination of diverse variables and complex inter-column relationships makes both datasets and particularly suitable for evaluating how well generative models preserve inter-column relationships in real-world tabular data. Moreover, these datasets presents a significant challenge for existing tabular data generation methods, which are primarily designed to enhance machine learning efficiency rather capture inter-column relationships.
4.2 Baselines
We compare our proposed method, LLM-TabFlow, against five representative approaches, including:
  • SMOTE (Chawla et al., 2002), an interpolation-based method originally introduced in the early 2000s to address dataset imbalance, which continues to outperform many modern generative models (Margeloiu et al., 2024).
  • CTGAN (Xu et al., 2019), a widely used GAN-based method known for its strong performance and industrial relevance.
  • TabDDPM (Kotelnikov et al., 2023) and TabSyn (Zhang et al., 2023), two diffusion model-based approaches leveraging iterative noise refinement for enhanced data synthesis—TabDDPM being the first of its kind and TabSyn representing the current state-of-the-art in this category.
  • GReaT (Borisov et al., 2022), an LLM-based method that reformulates tabular data as textual sequences to improve representation learning.
These methods span multiple generative paradigms, providing a diverse benchmark to evaluate how LLM-TabFlow performs relative to state-of-the-art techniques across different modeling approaches.
4.3 Synthetic Data Evaluation Framework
Figure 3: Real-world synthetic tabular data evaluation framework.
We introduce an evaluation framework to assess the performance of tabular data generation methods. As illustrated in Figure 3, the framework consists of three key assessment dimensions: data fidelity, data utility, and data privacy, each includes multiple sub-dimensions evaluated through a set of performance metrics. Data fidelity serves as the foundation, assessing whether the generated data accurately preserves the statistical properties, diversity, and inter-column relationships of real data. This is essential for ensuring the reliability and practical applicability of synthetic data. Building on fidelity, data utility examines the effectiveness of synthetic data in downstream tasks, including both machine learning tasks and broader real-world use cases. This dimension ensures the generated data is effective in real-world scenarios, such as business analytics, industrial simulations, and decision-making. In addition to data fidelity and utility, privacy is a crucial dimension in evaluating synthetic data generation. It measures the extent to which sensitive information is retained in the synthetic data, ensuring it does not compromise confidentiality or pose re-identification risks. Metrics like Distance to Closest Record (DCR) can be adopted to quantify privacy preservation. Together, these three dimensions would quantify whether synthetic data is statistically accurate, practically useful, and privacy-preserving—the essential aspects that synthetic data must be satisfied in real-world industrial use cases. We detail these three dimensions in the following sections.
4.3.1 Data Fidelity
Data fidelity includes three sub-dimensions: accuracy, diversity, and inter-column relationship, as detailed in the following sections.
Accuracy
Accuracy evaluates the degree to which synthetic data replicate the characteristics and patterns of real data, considering both low- and high-order statistics. We use column-wise distribution density estimation and pairwise column correlation for measuring low-order statistics, as described in Zhang et al. (2023). Specifically, Kolmogorov-Smirnov Test (KS test) and Total Variation Distance are employed to estimate column-wise density for continuous and categorical columns, respectively. Pearson correlation and contingency similarity are used to calculate pairwise column correlation for continuous and categorical columns, respectively. For accuracy in terms of high-order statistics, we employ $\alpha$-Precision (Alaa et al., 2022), which quantifies the local similarity between synthetic and real data points. This metric ensures that the synthetic data accurately represents the feature space of the real data, preserving fine-grained statistical properties, which is particularly crucial for industrial applications. Additionally, we use the Classifier Two Sample Test (C2ST) (Zein and Urvoy, 2022) to the similarity between synthetic and real data. Higher C2ST scores indicate a greater similarity to real data, making it more realistic to a classifier.
Diversity
Beyond accurately capturing data distributions, diversity plays a crucial role in evaluating synthetic data quality. This is particularly vital in industrial applications, where synthetic data must preserve all relevant variations and categories from the real dataset without omitting critical aspects. Both low- and high-order statistics are employed for evaluating diversity. It is crucial to ensure that all possible categories and value ranges from the real data are sufficiently represented in the synthetic data. To quantify this, we adopt average coverage score, a low-order statistics that evaluates how comprehensively a given region or feature space is covered by the synthetic data. It is mathematically defined as:
$C_{j}=\begin{cases}\frac{1}{N}\sum\limits_{j=1}^{N}\mathbb{1}\left[\min(X_{i})\leq X^{{}^{\prime}}_{i}(j)\leq\max(X_{i})\right]&\text{if }X^{{}^{\prime}}_{i}\text{ is continuous},\\ \frac{1}{N}\sum\limits_{j=1}^{N}\mathbb{1}\left[X^{{}^{\prime}}_{i}(j)\in X_{i}\right]&\text{if }X^{{}^{\prime}}_{i}\text{ is categorical}.\end{cases}$ (9)
[…]
We also incorporate $\beta$-Recall (Alaa et al., 2022), a high-order statistics, to evaluate diversity. This metric ensures that edge cases and less common combinations of features are adequately represented in the synthetic data. Such coverage is crucial for generating a more robust synthetic dataset, one that captures the full spectrum of variations present in real-world data, including uncommon and edge-case scenarios that might otherwise be overlooked.
Inter-Column Relationship
Finally, we evaluate inter-column logical relationship preservation (Long et al., 2025b), which in turn consists of two key dimensions: consistency and dependency. These aspects are critical in determining the quality of synthetic tabular data. The evaluation is based on three metrics:
  • Hierarchical Consistency Score (HCS): Measures the preservation of hierarchical relationships, such as city-to-country mappings, by quantifying the proportion of valid attribute groupings.
  • Multivariate Dependency Index (MDI): Assesses whether synthetic data maintains predefined functional dependencies, ensuring logical consistency across related attributes.
  • Distributional Similarity Index (DSI): Captures implicit dependencies by comparing the log-likelihoods of Gaussian Mixture Models fitted to real and synthetic datasets, identifying subtle structural deviations.
These dimensions collectively ensure that synthetic data retains key structural relationships, making it suitable for downstream applications in industrial contexts.
4.3.2 Data Utility
The utility of synthetic data is crucial in real-world industrial applications, where it can be applied to machine learning taskssuch as machine failure prediction and delivery delay prediction, both of which require high levels of accuracy for operational success. As shown in previous studies, Machine Learning Efficiency is a key factor in evaluating data utility.
To assess the effectiveness of synthetic data in machine learning tasks, we follow an evaluation process widely adopted in prior works (Lee et al., 2023; Kotelnikov et al., 2023; Zhang et al., 2023). First, generative models are trained on the real dataset to produce synthetic datasets of the same size as the real dataset. Machine learning models, such as an XGBoost-based classifier or regressor (Chen and Guestrin, 2016), are then trained on the synthetic data but evaluated on real dataset to measure generalization performance. For classification tasks, we evaluate performance using metrics like the Area Under the Receiver Operating Characteristic Curve (AUC) and the F1 score. For regression tasks, we assess predictive accuracy using the Coefficient of Determination ($R^{2}$), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE). Evaluating data utility is essential as it directly reflects the practical value of synthetic data. High test accuracy, when the model is trained on synthetic data, indicates that the synthetic dataset effectively preserves the underlying patterns and structures of the real data. This evaluation is crucial for validating the applicability of synthetic data in real-world industrial machine learning tasks.
Moreover, many industrial applications—such as operational decision-making, robust data-driven modeling, and scenario simulations—require synthetic data to be sufficiently realistic in mimicking real data to ensure reliable and actionable insights. Therefore, evaluating the utility of synthetic data must go beyond machine efficiency to its applicability in real-world scenarios. This involves evaluating the logical relationships within the data to ensure alignment with real-world operational dynamics, such as constructing retail networks, adhering to domain-specific rules, or simulating transportation events. By measuring how well the synthetic data meets these criteria, we can determine its reliability in driving actionable insights and realistic simulations.
4.3.3 Data Privacy
Synthetic data must preserve privacy by preventing the unintended disclosure of sensitive information from real-world industrial datasets and mitigating risks such as data re-identification and membership inference attacks. Therefore, assessing data privacy is a crucial dimension in evaluating the quality and reliability of synthetic data. To quantify privacy preservation, we employ the Distance to Closest Record (DCR) metric (Liu et al., 2024), which measures the pairwise distance between real and synthetic data points. This metric determines whether the synthetic data is generated based on the overall distribution of the real data, rather than replicating specific records from the training dataset. The DCR between each synthetic data point $s_{i}\in S$ and all real data points $r_{j}\in R$ is calculated using the $L_{1}$ (Manhattan) distance, which is the sum of the absolute differences between corresponding features. For each synthetic data point $s_{i}$, the minimum distance to any real data point $r_{j}$ is computed and denoted as $d_{\text{min}}(s_{i},R)$, representing the closest real data point to $s_{i}$. Similarly, the minimum distance between each synthetic data point $s_{i}$ and all test data points $t_{k}\in T$ is computed in the same manner, with the minimum distance denoted as $d_{\text{min}}(s_{i},T)$, representing the closest test data point to $s_{i}$. The DCR is then defined as the proportion of synthetic data points whose minimum distance to the real data is smaller than the minimum distance to the test data:
$\text{DCR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}\Bigl{(}d_{\text{min}}(s_{i},R)<d_{\text{min}}(s_{i},T)\Bigr{)}$ (10)
where $N$ is the total number of synthetic data points, and $\mathbb{1}(\cdot)$ is an indicator function that returns 1 if the condition is satisfied and 0 otherwise.
This DCR score quantifies how closely the synthetic data points resemble the real data, with a higher score indicating that the synthetic data is closer to real data, which could imply a privacy risk. A lower DCR score suggests that the synthetic data points are more distinct from the real data, thereby better preserving privacy.
4.4 Experimental Setup
Data Preparation
The two evaluation datasets were first split into training and testing sets, following the separation ratio specified in Table 1. Tabular data generation methods were then trained on the real training set, and the trained models were used to generate synthetic dataset with the same size as the real training set.
Evaluation
In our experiments, we employed GPT-4 (Achiam et al., 2023) for column reasoning, enabling LLM-TabFlow to capture complex inter-column relationships effectively. To ensure a robust evaluation, we generated ten synthetic datasets using each trained tabular generation method. Each synthetic dataset was then used to train task-specific models—specifically, an XGBoost classifier in our experiments—for downstream tasks such as prediction. We then calculated the mean and standard deviation of the results.
Hyperparameters
The hyperparameters of the prediction model—XGBoost classifier—were search within the following parameter space: number of trees: {100, 200, 300}; learning rate: {0.1, 0.01, 0.001}; maximum tree depth: {5, 10}; minimum child weight: {1, 10}; and gamma: {0.0, 0.5, 1.0}. Additionally, Grid Search was employed to identify the optimal hyperparameter combination for the XGBoost classifer.
Software and Hardware
We implemented all the experiments in Python. The implementations of the tabular data generation methods and the XGBoost classifier were sourced from their original papers. For evaluation metrics, we used those available in SDMetrics  [^2] https://docs.sdv.dev/sdmetrics; otherwise, we implemented them ourselves. Our experiments were conducted in the workstation with the following configurations: Ubuntu 20.04.6 LTS with the Linux kernel version 5.15.0-113-generic, an Intel(R) Xeon(R) Platinum 8368 CPU (2.40 GHz), and an NVIDIA GeForce RTX 4090 GPU. These configurations are well-suited for high-performance computing and machine learning tasks.
5 Experiments Results
(a) Retailing
(b) Purchasing
Figure 4: Radar Charts illustrating the performance of all methods on the two datasets across the six dimensions: Accuracy, Diversity, Utility, Privacy, Consistency, and Dependency. Values closer to the outermost circle indicate better performance. LLM-TabFlow achieves the overall best performance, with the most balanced results across all metrics.
This section presents the evaluation results for synthetic data produced by our proposed method—LLM-TabFlow—and baseline methods. Each evaluation is conducted ten times, repeating the process of generating and evaluating synthetic data. The results are reported as the results presented as the mean and standard deviation in the format ${\tt mean}\pm{\tt stddev}$.
5.1 Overall Evaluation
The overall performance of LLM-TabFlow and the five baseline methods on the two datasets is visualized as Radar charts, as illustrated in Figure 4. To provide a more comprehensive evaluation and clearer visualization, the inter-column relationship dimension is divided into two distinct aspects: Consistency and Dependency. As a result, the radar charts consist of six dimensions: Accuracy, Diversity, Utility, Privacy, Consistency, and Dependency. Surprisingly, as shown in Figure 4, SMOTE, originally proposed in early 2000s, still outperforms advanced deep generative models in several aspects such as accuracy, diversity, and inter-column relationships, while also achieving strong performance in utility on the Purchasing dataset. However, compared to deep generative models such as latent-space or PLM-based methods, this interpolation-based approach heavily relies on real input data for data sampling and often compromising data privacy. These limitations significantly restrict its applicability in industrial scenarios involving sensitive data. In contrast, advanced models such as TabDDPM and CTGAN perform slightly worse than SMOTE across almost all metrics. While TabSyn, a state-of-the-art method, surpasses SMOTE in privacy and utility, it struggles to fully capture consistency and dependency in the real data. Overall, models relying on latent representation exhibit significant limitations in preserving logical relationships. Notably, GReaT, a PLM-based model, excels in maintaining strong logical relationships but falls short in other critical dimensions, including data fidelity and utility. In contrast, our proposed approach—LLM-TabFlow—achieves near-perfect results across all six dimensions, ensuring consistency and dependency while excelling in fidelity, utility, and privacy. This demonstrates the robustness and versatility of LLM-TabFlow in generating high-quality realistic synthetic tabular data. We further examine the experimental results in each of the six dimensions in the following sections.
5.2 Accuracy and Diversity
Table 2: Evaluation results of the six methods in terms of accuracy (density estimation, Pearson correlation, and $\alpha$-Precision) and diversity (average coverage score and $\beta$-Recall). Higher values of these metrics indicate better performance. The best and second-best results are highlighted in blue and black boldfaced, respectively.
Datasets Metrics Interpolation-based Latent Space-based PLM-based Ours
SMOTE CTGAN TabDDPM TabSyn GReaT LLM-TabFlow
Retailing Density Estimation 98.02±0.02 90.38±0.03 33.11±0.02 96.38±0.04 89.58±0.02 96.46±0.03
Pairwise Correlation 96.21±0.62 74.41±0.14 36.78±0.01 94.81±0.04 71.00±1.16 92.52±0.42
$\alpha$-Precision 93.54±0.01 88.90±0.14 0.00±0.00 98.48±0.10 82.18±0.15 99.15±0.12
Average Coverage Score 99.41±0.08 82.27±0.14 76.23±0.23 99.52±0.11 92.34±0.15 99.18±0.14
$\beta$-Recall 72.21±0.13 1.71±0.01 0.00±0.00 22.62±0.12 24.05±0.14 27.85±0.13
Purchasing Density Estimation 99.11±0.05 86.36±0.04 48.01±0.04 98.61±0.05 92.72±0.08 98.14±0.07
Pairwise Correlation 98.83±0.11 82.25±0.33 51.00±0.04 98.63±0.20 74.18±3.32 92.34±1.12
$\alpha$-Precision 98.01±0.32 73.67±0.22 4.16±0.01 96.73±0.37 88.00±0.36 97.29±0.25
Average Coverage Score 93.19±2.34 75.87±0.13 97.49±0.80 91.85±0.45 88.72±2.97 93.92±0.23
$\beta$-Recall 62.79±0.63 4.23±0.08 0.03±0.04 15.30±0.37 36.16±0.32 17.32±0.21
(a) Item Profit Ratio
(b) Product Price
(c) Latitude
Figure 5: Density plots for the three continuous columns (item profit ratio, product price, and latitude), comparing the distribution of real data and their synthetic counterparts generated by different methods. Curves that more closely align with the real data indicate better performance. Both LLM-TabFlow and TabSyn exhibit distributions that closely match the real data, outperforming other methods.
(a) Shipping Mode
(b) Payment Type
(c) Order Status
Figure 6: Distribution plots for the three categorical columns (shipping mode, payment type, and order status), comparing synthetic data to real data. Distributions that closely match the real data indicate superior performance. Both LLM-TabFlow and TabSyn exhibit distributions that are significantly closer to the real data compared to other methods.
(a) SMOTE
(b) CTGAN
(c) TabDDPM
(d) GReaT
(e) TabSyn
(f) Ours (LLM-TabFlow)
Figure 7: Heatmap illustrating the absolute divergence in pairwise column correlations between the synthetic and real data. Lighter colors indicate smaller differences and better alignment. SMOTE, TabSyn, and LLM-TabFlow exhibit the closest alignment with the real data, outperforming other methods.
(a) SMOTE
(b) CTGAN
(c) TabDDPM
(d) GReaT
(e) TabSyn
(f) Ours (LLM-TabFlow)
Figure 8: Visualization of the coverage score for both categorical and numerical columns using Radar charts. Labels starting with N and C represent numerical and continues columns around circles, respectively. LLM-TabFlow, TabSyn, and SMOTE significantly outperform the other three methods in terms of coverage across all evaluated columns.
Data fidelity is evaluated using a set of data accuracy and diversity metrics. The main objective of data fidelity evaluation is to assess whether the synthetic data is sufficiently authentic and accurate when compared to the real data.
For data accuracy, we first utilize low-order statistics, such as column-wise density estimation and pairwise correlation scores (i.e., Pearson Correlation), to measure the general statistical properties of the columns. As shown in Table 2, SMOTE achieves the overall best performance on both datasets, followed closely by LLM-TabFlow and then TabSyn. Secondly, for high-order statistics, which capture more complex statistical relationships, the $\alpha$-Precision scores demonstrate that LLM-TabFlow outperforms all other methods, achieving 99.15% and 93.92% on the Retailing and Purchasing datasets, respectively. TabSyn and SMOTE follow closely, with values of 98.48% and 93.19%. To further examine the accuracy of synthetic data generated by various models, we analyzed the density distributions of representative columns in the Retailing dataset. First, we presneted the distribution of the three continuous columns—item profit ratio, product price, and latitude—in Figure 5. As shown in Figure 5, TabSyn and LLM-TabFlow perform the overall best, producing distributions that are nearly identical to the real data. While CTGAN and GReaT outperform than TabDDPM and generate distribution somewhat close to the real data, they exhibit noticeable gaps in peak regions, showing biases to the real data. SMOTE presents lower variations in these regions, but it is overall slightly less accurate than TabSyn and LLM-TabFlow. For categorical columns, such as shipping mode, payment type, and order status, as shown in Figure 6, LLM-TabFlow outperforms all other methods, accurately replicating category counts in these columns. TabSyn performs slightly worse than LLM-TabFlow but better than SMOTE. These results highlight LLM-TabFlow and TabSyn is superior to SMOTE for categorical data generation. Additionally, we visualize the absolute divergence in pairwise column correlations between the real Retailing dataset and its synthetic counterpart using heatmaps, providing a clear and intuitive representation of the differences. As presented in Figure 7, SMOTE, Tabyn, and LLM-TabFlow, though comparable to each other, achieve much better performance than other methods (e.g., CTGAN, TabDDPM, and GReaT) in terms of pairwise column correlation, with a closer match to the real data.
For the data diversity, we continue by evaluating how well the synthetic data captures all modes present in the real data using the average coverage score described in Equation 9. The experimental results are presented in Table 2. As shown in Table 2, for single columns, LLM-TabFlow and SMOTE achieve the highest values on the Purchasing and Retailing datasets, respectively. In addition, we illustrate the coverage scores of synthetic data generated by each method for categorical and numerical columns in the Retailing dataset using radar charts. As illustrated in Figure 8, CTGAN excels in covering more modes for categorical data, while TabDDPM captures a broader range of numerical data. GReaT performs comparably for both numerical and categorical columns but fails to cover numerical column 7 (See 8(d)). LLM-TabFlow, Tabsyn, and SMOTE nearly cover all modes for both categorical and numerical columns. To assess the diversity of joint feature interactions, we used $\beta$-Recall scores. SMOTE, which explicitly balances minority classes, achieves the highest scores across both datasets, followed by LLM-TabFlow and GReaT, which rank second for each dataset, respectively. Notably, TabDDPM performs poorly on both accuracy and diversity metrics, indicating its inability to generate sufficiently diverse values. In summary, LLM-TabFlow strikes an excellent balance between accuracy and diversity, outperforming other models in most aspects, particularly in synthesizing continuous and categorical data.
5.3 Inter-Column Relationship Preservation
Table 3: Inter-column relationship identified in the Retail and Purchasing datasets.
Dataset Relationship Group Attributes Condition
Retailing Hierarchical Consistency Geographical (Orders)
$x_{1,j}$ (order city), $x_{2,j}$ (order state), $x_{3,j}$ (order country),
$x_{4,j}$ (order region), $x_{5,j}$ (order market)
$(x_{i,j})_{i\in G_{1}}\in C_{1,j}$
Geographical (Customers)
$x_{6,j}$ (customer city), $x_{7,j}$ (customer state),
$x_{8,j}$ (customer country)
$(x_{i,j})_{i\in G_{2}}\in C_{2,j}$
Products Information
$x_{9,j}$ (category ID), $x_{10,j}$ (category name), $x_{11,j}$ (department ID),
$x_{12,j}$ (department name), $x_{13,j}$ (product card ID),
$x_{14,j}$ (product category ID), $x_{15,j}$ (product name)
$(x_{i,j})_{i\in G_{3}}\in C_{3,j}$
Temporal
Dependency
Delivery Time $x_{1,j}$ (order date), $x_{2,j}$ (delivery time) $D_{1,j}:x_{1,j}<x_{2,j}$
Mathematical Dependency Sales Values $x_{1,j}$ (quantity), $x_{2,j}$ (price), $x_{3,j}$ (discount rate), $x_{4,j}$ (discount value), $x_{5,j}$ (original price), $x_{6,j}$ (sales price) $D_{2,j}:x_{5,j}=x_{1,j}\times x_{2,j}$
$D_{3,j}:x_{5,j}=x_{1,j}\times x_{2,j}\times x_{3,j}$
$D_{4,j}:x_{4,j}=x_{7,j}-x_{5,j}$
Purchasing
Hierarchical
Consistency
Supplier Information $x_{1,j}$ (supplier number), $x_{2,j}$ (supplier name) $(x_{i,j})_{i\in G_{1}}\in C_{1,j}$
Temporal Dependency Delivery Time $x_{3,j}$ (order created date), $x_{4,j}$ (planned delivery time), $x_{5,j}$ (realistic delivery time) $D_{1,j}:x_{3,j}<x_{4,j}$
$D_{2,j}:x_{3,j}<x_{5,j}$
Mathematical Dependency Purchase Values $x_{6,j}$ (quantity), $x_{7,j}$ (price), $x_{8,j}$ (discount rate), $x_{9,j}$ (net amount), $x_{10,j}$ (gross amount) $D_{3,j}:x_{9,j}=x_{6,j}\times x_{7,j}$
$D_{4,j}:x_{10,j}=x_{9,j}\times(1-x_{8,j})$
Table 4: Evaluation results for preserving inter-column relationships in synthetic tabular data. Higher values indicate better performance. The best and second-best results are highlighted in blue and black boldfaced, respectively.
Datasets Metrics Interpolation Latent Space-based PLM-based Ours
SMOTE CTGAN TabDDPM TabSyn GReaT LLM-TabFlow
Retailing HCS 98.09$\pm$0.01 39.23$\pm$0.03 16.07$\pm$0.01 71.63$\pm$0.08 98.01$\pm$0.01 100.00$\pm$0.00
MDI 87.03$\pm$0.03 38.87$\pm$0.05 59.08$\pm$0.05 68.34$\pm$0.08 97.37$\pm$0.02 100.00$\pm$0.00
DSI 74.94$\pm$0.00 30.55$\pm$0.01 79.72$\pm$0.02 66.94$\pm$0.01 98.35$\pm$0.01 98.36$\pm$0.01
Purchasing HCS 81.17$\pm$0.29 0.00$\pm$0.00 0.01$\pm$0.03 54.99$\pm$0.23 99.92$\pm$0.01 100.00$\pm$0.00
MDI 84.82$\pm$0.09 36.69$\pm$0.06 63.49$\pm$0.13 56.30$\pm$0.11 94.54$\pm$0.06 100.00$\pm$0.00
DSI 99.75$\pm$0.02 59.39$\pm$0.01 96.62$\pm$0.01 93.15$\pm$0.00 94.45$\pm$0.02 96.83$\pm$0.02
Through LLM reasoning, the inter-column relationships in the two datasets are identified, as presented in Table 3. These relationships are categorized into different types, with corresponding attributes and conditions specified for both datasets. To quantify the inter-column consistency and dependency, we employ three evaluation metrics—HCS, MDI, and DSI—proposed in Long et al. (2025b). The results are presented in Table 4. As shown, LLM-TabFlow achieves a perfect value of 100% in both consistency and dependency, demonstrating its ability to fully preserve the logical relationships within the real data. GReaT performs slightly worse, exhibiting strong inter-column relationship preservation but with some errors in capturing dependencies across multiple columns. In contrast, latent space-based generative models (e.g., CTGAN, TabDDPM, and TabSyn) perform significantly worse, particularly in maintaining hierarchical consistency across both datasets. Interestingly, SMOTE outperforms most advanced generative models in preserving inter-column logical relationships. To further assess dependency preservation, we evaluate the DSI score by analyzing previously unseen relationships between variables and target values. Specifically, we examine financial attributes to assess their impact on target outcomes across both datasets. LLM-TabFlow achieves the overall best performance, while SMOTE and GReaT perform well on individual datasets. These results highlight LLM-TabFlow’s strength in preserving both simple logical relationships and complex inter-column dependencies within tabular data, setting a new benchmark for synthetic tabular data generation with preserved inter-column relationships.
5.4 Utility: Machine Learning Efficiency
Table 5: Evaluation results of machine learning (classification) efficiency in terms of the AUC score. Higher AUC values indicate better performance. The best and second-best results are highlighted in blue and black boldfaced, respectively.
Real Dataset Interpolation Latent Space-based PLM-based
Ours
Datasets Real SMOTE CTGAN TabDDPM TabSyn GReaT LLM-TabFlow
Retailing 88.17 69.59±0.33 68.91±0.18 49.68±0.91 71.55±0.24 71.15±0.14 71.90±0.20
Purchasing 84.37 83.96±0.44 53.31±0.60 49.72±1.28 62.26±1.38 50.06±0.75 63.45±0.92
While data utility, as illustrated in Figure 3, consists of two sub-dimensions for evaluation, this experiment mainly focuses on quantitatively assessing the utility of the generated synthetic data through its machine learning efficiency. The experimental results in Table 5 demonstrate that LLM-TabFlow achieves the overall best performance across both datasets, obtaining the highest AUC on the Retailing dataset and the second-highest on the Purchasing dataset. Specifically, in the Retailing dataset, our approach outperforms TabSyn in terms of AUC scores for classification tasks. GReaT is just behind TabSyn but surpasses SMOTE, demonstrating the advantages of LLM-based training and sampling in generating synthetic tabular data. In contrast, GANs and TabDDPM models are considerably less effective in improving classification efficiency with large-scale synthetic data. However, all these methods still lag significantly behind those trained on the real dataset. In the Purchasing dataset, SMOTE achieves highest AUC score and achieves performance close to that of real data. LLM-TabFlow, which achieves the second-best AUC, along with other methods, still lags behind real data, highlighting the ongoing challenge of achieving parity with real data in machine learning efficency, despite the advancements in generative modeling techniques. Moreover, the ability of synthetic data to preserve inter-column relationships directly impacts its utility in real-world applications, such as process simulation and decision-making in finance and delivery of supply chains. Based on the results in Table 4 and Table 5, LLM-TabFlow generates synthetic tabular data that can effectively substitute real data, achieving satisfactory performance in real-world machine learning tasks.
5.5 Privacy Preservation
Table 6: Evaluation results of privacy preservation in terms of the DCR and C2ST. Lower DCR values indicate better privacy preservation, with the best and smallest value being 50%. Higher C2ST values indicate better performance in terms of accuracy. The best and second-best results are highlighted in bold and black boldfaced, respectively.
Datasets Metrics Interpolation Latent Space-based PLM-based Ours
SMOTE CTGAN TabDDPM TabSyn GReaT LLM-TabFlow
Retailing DCR 99.85±0.01 89.80±0.06 89.15±0.08 90.18±0.07 90.19±0.05 90.03±0.03
C2ST 92.64±0.29 17.65±0.32 0.00±0.00 52.89±0.23 38.83±0.06 66.75±0.15
Purchasing DCR 96.62±0.15 90.67±0.20 95.02±0.21 90.60±0.21 91.89±0.25 90.43±0.22
C2ST 98.11±0.57 66.92±0.69 0.27±0.13 93.93±0.39 76.34±0.37 95.38±0.30
We further evaluate the proposed method alongside the five baseline approaches in the privacy dimension. As described in Section 5.5, DCR (see Equation 10)—a metric that quantifies the similarity between synthetic and real data—is used to evaluate privacy preservation. Additionally, since balancing privacy and utility is crucial for effective tabular data generation, we adopt the Classifier Two-Sample Test (C2ST) (see Section 4.3.1) to measure the distinguishability of real and synthetic data. For this experiment, a logistic regression classifier is used to compute the C2ST. The results in Table 6 show that LLM-TabFlow and TabDDPM achieve the best privacy-preserving performance (highest DCR values) on the Purchasing and Retailing datasets, respectively. However, TabDDPM obtains an extremely low C2ST value, indicating that its strong privacy preservation performance comes at the expensive cost of accuracy. In contrast, SMOTE and LLM-TabFlow achieve the highest and second-highest C2ST scores on the Retailing and Purchasing dataset, respectively, demonstrating strong accuracy in data synthesis. While SMOTE shows the best accuracy in terms of accuracy (highest C2ST value), it achieves the worst (lowest DCR value) privacy preservation performance, highlighting its high accuracy cost at preserving privacy. Our method, however, provides the best balance between privacy and accuracy, outperforming both SMOTE and state-of-the-art generative models. This balance is crucial for applications that require both robust data privacy protection and high data utility, making our approach particularly well-suited for real-world industrial scenarios that demand both privacy guarantees and accurate data synthesis.
6 Discussion and Implications
We propose a novel approach—LLM-TabFlow—that integrates LLM reasoning with score-based diffusion models to generate synthetic tabular data. This approach significantly outperforms traditional diffusion-based models in capturing complex inter-column logical relationships. Unlike conventional methods that embed all features with intricate relationships directly into latent space, LLM-TabFlow introduces an innovative framework that compresses complex tabular data into simplified representations for generation. These representations are subsequently decompressed back into their original form after sampling. This approach not only enhances the model’s ability to handle complex relationships but also provides a more efficient and interpretable framework for tabular data synthesis. Experimental results demonstrate that PLM-based models excel at capturing inter-column relationships, which may due to their autoregressive nature. However, these methods struggle to accurately preserve the low- and high-order statistics of the original (real) dataset and require approximately ten times more training and inference time compared to latent space-based models. Additionally, it was observed that GReaT can generate unseen categories for product names, revealing the inherent uncertainty of LLM-based generative models. This finding underscores the need for caution when using such models to generate complex categorical data. In contrast, our approach mitigates the uncertainties, providing a more reliable method for tabular data generation. It outperforms existing latent-space-based models in accurately handling datasets with diverse logical dependencies. Furthermore, LLM-TabFlow enhances computational efficiency and improves the capture of heterogeneous data distributions in latent space compared to other LLM-based models. Overall, LLM-TabFlow achieves a much better balance between accuracy, utility, and privacy, positioning it as a robust solution for synthetic tabular data generation.
However, this work has some limitations. Firstly, the identification of relationships in the data is influenced by the knowledge of the LLMs. There is no guarantee that LLMs can capture all complex relationships, particularly when column names are abbreviated or the descriptions are ambiguous or inaccurate. This reliance on LLMs may introduce biases or variances in understanding the underlying data structure. Secondly, the experimental evaluation is limited to datasets within the supply chain domain. To ensure broader applicability, our method needs to be tested on more diverse and comprehensive tabular datasets from other industrial sectors, such as healthcare, finance, and social sciences. This would help verify the generalizaubility of LLM-TabFlow across different types of domain knowledge and data structures. Additionally, privacy concerns arise if column names and descriptions contain sensitive information, as LLMs process these inputs to infer relationships. This could potentially expose private information, especially in domains such as healthcare or finance.
Despite these limitations, LLM-TabFlow shows significant potential for real-world uses across various industries. Beyond preserving data privacy and enhancing machine learning tasks, it is, to the best of our knowledge, the first framework capable of capturing and replicating complex real-world dependencies in synthetic tabular data. This capability is particularly valuable for conducting simulations in digital twins (Xu et al., 2025), enabling more accurate business analysis, predictive modeling, and data-driven decision-making. For example, LLM-TabFlow can be applied to model complex relational data across multiple competing business entities, addressing the data scarcity problems in industrial domains caused by privacy and reluctance to share data (Xu et al., 2025; AlMahri et al., 2024). In agent-based simulations, LLM-TabFlow would generates synthetic data that accurately reflects the interactions between agents, enabling researchers to study complex systems such as supply chains (Xu et al., 2024b, 2025), financial markets (Proselkov et al., 2024), and social networks. Similarly, high-quality, realistic synthetic data would be used to train and fine-tune autonomous agents for customized scenarios, enhancing their capabilities in dynamic environments, such as managing supply chain operations and making sophisticated decision making in autonomous supply chains (Xu et al., 2024a, b, c). These industrial implications highlight LLM-TabFlow’s significant impact in generating high-fidelity, privacy-preserved synthetic data, positioning it as a transformative tool for driving data-driven innovation in both industry and academia.
7 Conclusion and Future Work
This paper addresses the challenge of preserving inter-column relationships in synthetic tabular data generation for real-world applications. We propose an evaluation framework comprising six dimensions to evaluate the performance of synthetic tabular data in industrial contexts, where both high utility and strong privacy preservation are essential. We have proposed LLM-TabFlow, a novel synthetic tabular data generation approach that integrates LLM reasoning with latent space-based data generation. Additionally, we conduct extensive experiments to evaluate the performance of the proposed method and the other five baselines under the proposed evaluation framework. These experiments show that existing methods struggle to preserve logical inter-column relationships, often failing to maintain dependencies and consistenc. In contrast, effectively upholds hierarchical consistency and logical dependencies while achieving the best trade-off among data fidelity, privacy, and utility.
To the best of our knowledge, LLM-TabFlow is the first solution capable of synthesizing tabular data while simultaneously preserving inter-column relationships and ensuring privacy. This highlights its potential for generating high-quality data for industrial applications, including simulation and advanced decision-making. Moreover, this approach lays the foundation for integrating LLM reasoning with tabular data generation, offering a pipeline that can be extended to other domains such as finance and healthcare. By doing so, LLM-TabFlow unlocks new opportunities for both research and real-world applications, fostering innovation in data-driven industries.
However, our study has certain limitations, as discussed in Section 6. Addressing these limitations will be a key focus of our future work. First, we will evaluate LLM-TabFlow on a broader range of diverse and challenging datasets, including those from various industrial domains and datasets with abbreviated or ambiguous column names and descriptions. This will help assess the model’s adaptability and its ability to capture complex inter-column relationships across various contexts. Second, we will explore to incorporate structural constraints to achieve more precise and context-aware data synthesis. Another key focus of our future work is evaluating LLM-TabFlow’s ability to identify inter-column relationships when sensitive information in column names or descriptions is anonymized. Furthermore, while LLM-TabFlow effectively preserves logical relationships in complex tabular data, real-world tables often contain domain-specific constraints that are not always apparent, such as environmental limits (e.g., industrial temperature ranges) and natural rules (e.g., age must be over 18 for adulthood). Incorporating new prompting methods into our model is crucial for maintaining these structured constraints, ensuring more precise and context-aware data synthesis. Additionally, preserving column relationships across multiple tables is essential for generating realistic synthetic tabular data, especially in scenarios involving foreign key dependencies. Our future work will focus on identifying and maintaining these cross-table relationships to ensure consistency, which is critical for information systems in industries such as supply chain management, e-commerce, and healthcare.